OpenStudy (anonymous):

Please help. Write a rule for the nth term of the sequence. Use your rule to find a100.–8, 9, 26, 43, 60, . . .

OpenStudy (anonymous):

This is an arithmetic progression with a distance of 17...So, each \[a_{k}\] can be calculated by \[a_{k-1}+17\] Now, \[a_{n}=a_{1}+(n-1)d\] And we want the n=100 from the progression So, you'll have \[a_{100}=-8+(100-1) \times 17\] \[a_{100}=1675\]